OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

jimthompson5910 (jim_thompson5910):

first you need to determine the slope so use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

i got 2

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now use the point slope form y-y1 = m(x-x1)

OpenStudy (anonymous):

so far i got y-0=2x-4

jimthompson5910 (jim_thompson5910):

you keep it in point slope form

jimthompson5910 (jim_thompson5910):

so don't distribute

OpenStudy (anonymous):

so y-0=2(x-{-2})

OpenStudy (anonymous):

I got it! thanks.

jimthompson5910 (jim_thompson5910):

good, you can simplify things to get y-0 = 2(x+2)